# Linear Discriminant Regularized Regression

### v1

@@ 2 Methodology
This, in turn, requires that we estimate the discriminant directional matrix $B^{*}=\Sigma_{w}^{-1}{\mathbb{M}}$. The following lemma establishes the connection between $B^{*}$ and the regression matrix $B$ in (1.13). Recall the definitions of $\pi$, $D_{\pi}$ and ${\mathbb{M}}$ from (1.5) as well as $\Sigma={\rm Cov}(X)$.
@@
Under model (1.1) and (1.3), we have
@@
$\Sigma_{XY}=\mathbb{E}(XY^{\top})-\mathbb{E}(X)(\mathbb{E}(Y))^{\top}={\mathbb{M}}D_{\pi}$ (2.4)
@@
$\Sigma=\Sigma_{w}+\Sigma_{b}=\Sigma_{w}+{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top},$ (2.6)
@@
$\Sigma_{w}^{-1}=\Sigma^{-1}+\Sigma^{-1}{\mathbb{M}}\left(\mathbb{I}_{L}-D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}\right)^{-1}D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}.$
@@
$B^{*}$ $=\Sigma_{w}^{-1}{\mathbb{M}}$
$=\Sigma^{-1}{\mathbb{M}}+\Sigma^{-1}{\mathbb{M}}\left(\mathbb{I}_{L}-D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}\right)^{-1}D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}$
$=\Sigma^{-1}{\mathbb{M}}\left(\mathbb{I}_{L}-D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}\right)^{-1}$
$=\Sigma^{-1}{\mathbb{M}}D_{\pi}\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}D_{\pi}\right)^{-1}$
$=B\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B\right)^{-1}$
@@ 2.1 Estimation of the regression matrix
In view of (1.13), it is natural to estimate $B$ via regularized regression methods in (1.14). Choices of the penalty term $\text{pen}(B)$ depend on the concrete problem at hand. For example, if we believe that only a subset of the original features are important for predicting the label, then sparsity structures of the regression matrix $B$ are reasonable assumptions. If the subset of features, that are pertinent for predicting the label, varies across different levels, then the entry-wise lasso penalty is appropriate. To fix ideas and due to technical complications, we focus our analysis to the matrix estimator $\widehat{B}$ in (1.14) obtained with the following concrete lasso penalty
@@
for some tuning parameter $\lambda>0$. It is fairly straightforward to modify our analysis to other choices of the penalty term discussed below.
Remark 1 (Alternative penalty choices).
If there exist many features that are not predictive for all levels of the response label, then the group-lasso penalty $\text{pen}(B)=\lambda\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ is a more suitable choice. In this case, one can retain a much smaller subset of important features and achieve feature selection. Furthermore, one can also consider either a row-wise or a column-wise elastic net penalty if the features tend to (highly) correlate with each other. Another example that is commonly used in the literature is when entries of $B$ exhibit certain smoothness structure, such as some imaging or audio applications in Hastie et al. (1995). In this case, one could choose $\text{pen}(B)={\rm tr}(B^{\top}\Omega B)$ for some pre-specified “roughness” penalty matrix $\Omega$. On the other hand, if the goal is to construct a low-dimensional subset of linear combinations of the original feature for classification purpose, then a low-rank penalty in (1.14) could be more appropriate. For example, if the matrix of conditional means ${\mathbb{M}}\in\mathbb{R}^{p\times L}$ has a reduced rank $r\ll\min\{p,L\}$ such as under the model in Bing and Wegkamp (2023), then Lemma 1 implies that ${\rm rank}(B)\leq r$. Estimation of $B\in\mathbb{R}^{p\times L}$ in (1.14) could be done by using either the reduced rank penalty $\text{pen}(B)=\lambda~{}{\rm rank}(B)$ or the nuclear norm penalty $\text{pen}(B)=\lambda\|B\|_{*}$. We will investigate this approach in another paper.
@@ 2.2 Estimation of the discriminant matrix and the Bayes rule
The discriminant functions in (2.1) that form the basis of the Bayes classification rule (1.4) necessitate that we estimate the matrix of the conditional means ${\mathbb{M}}$, the vector of prior probabilities $\pi$ as well as the discriminant directional matrix $B^{*}$. We adopt the following standard estimators
$\widehat{\mathbb{M}}={\mathbb{X}}^{\top}{\mathbb{Y}}({\mathbb{Y}}^{\top}{\mathbb{Y}})^{-1},\qquad D_{\widehat{\pi}}={\rm diag}(\widehat{\pi})={1\over n}{\mathbb{Y}}^{\top}{\mathbb{Y}}$ (2.8)
to estimate ${\mathbb{M}}$ and $D_{\pi}$, respectively. Recall from Lemma 1 that $B^{*}=BH^{-1}$. Since the estimator $\widehat{B}$ of $B$ is computed in (1.14), in view of the closed-form of $H$ in (2.3), we propose to estimate $H$ by the plug-in estimator
$\widehat{H}=D_{\widehat{\pi}}-{1\over n}\widehat{B}^{\top}{\mathbb{X}}^{\top}{\mathbb{X}}\widehat{B}={1\over n}\left({\mathbb{Y}}^{\top}{\mathbb{Y}}-\widehat{B}^{\top}{\mathbb{X}}^{\top}{\mathbb{X}}\widehat{B}\right).$ (2.9)
Note that ${\mathbb{X}}\widehat{B}$ is simply the in-sample fit of the regression in (1.14). Computation of ${\mathbb{X}}\widehat{B}$ thus can be done efficiently even in non-parametric regression settings, for instance, when the columns of ${\mathbb{X}}$ consist of basis expansions. Now, using $B^{*}=BH^{-1}$, we can further estimate $B^{*}$ by
@@
Here $M^{+}$ denotes the moore-penrose inverse of any matrix $M$. Finally, in view of (2.1), we estimate the discriminant functions by
$\widehat{G}_{\ell}(x)=\widehat{\mu}_{\ell}^{\top}\widehat{B}^{*}_{\ell}-2x^{\top}\widehat{B}^{*}_{\ell}-2\log(\widehat{\pi}_{\ell})\quad\text{for all $\ell\in[L]$. }$ (2.11)
@@
Theoretically, we will show that $\widehat{H}$ is invertible with overwhelming probability for the lasso-penalty (see, Remark 3 as well as the proof of Theorem 3 in Section A.3). Moreover, it is worth mentioning that the inverse of $\widehat{H}$ always exists when any $\ell_{2}$-norm related penalty is deployed in (1.14).
@@ 2.3 Classification based on reduced dimensions
We start with the fact (its proof is stated in Section A.1 for completeness) that the Bayes rule in (1.4) is equivalent to
@@
$C_{b}:=B^{\top}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}B.$
For a given reduced dimension $1\leq K<L$, we can find $K$ discriminant directions in the space of $Z$ by solving, for each $k\in[K]$, the minimization problem
$\begin{split}\alpha_{k}=\operatorname*{arg\,max}_{\alpha\in\mathbb{R}^{L}}&~{}\alpha^{\top}C_{b}~{}\alpha\\ \text{such that }&~{}\alpha^{\top}C_{w}~{}\alpha=1,~{}\alpha^{\top}C_{w}~{}\alpha_{i}=0,~{}~{}\text{ for all }~{}i<k.\end{split}$
@@
$\widehat{C}_{b}=\widehat{B}^{\top}\widehat{\mathbb{M}}D_{\widehat{\pi}}\widehat{\mathbb{M}}^{\top}\widehat{B}={1\over n}\widehat{B}^{\top}{\mathbb{X}}^{\top}P_{{\mathbb{Y}}}{\mathbb{X}}\widehat{B},\qquad\widehat{C}_{w}={1\over n}\widehat{B}^{\top}{\mathbb{X}}^{\top}(\mathbb{I}_{n}-P_{{\mathbb{Y}}}){\mathbb{X}}\widehat{B},$
with $P_{{\mathbb{Y}}}={\mathbb{Y}}({\mathbb{Y}}^{\top}{\mathbb{Y}})^{-1}{\mathbb{Y}}^{\top}$. Subsequently, the estimated discriminant directions $\widehat{\alpha}_{k}$, for $1\leq k\leq K$, are computed by solving
$\begin{split}\widehat{\alpha}_{k}=\operatorname*{arg\,max}_{\alpha\in\mathbb{R}^{L}}&~{}\alpha^{\top}\widehat{C}_{b}~{}\alpha\\ \text{such that}&~{}\alpha^{\top}\widehat{C}_{w}~{}\alpha=1,\quad\alpha^{\top}\widehat{C}_{w}~{}\widehat{\alpha}_{i}=0,~{}~{}\forall~{}i<k,\end{split}$ (2.16)
@@
$\text{span}({\mathbb{X}}\widehat{B})\subset\text{span}({\mathbb{X}})$
@@
${\rm rank}((\mathbb{I}_{n}-P_{{\mathbb{Y}}}){\mathbb{X}}\widehat{B})={\rm rank}({\mathbb{X}}\widehat{B}),$ (2.17)
we do not suffer numerically in solving (2.16), in stark contrast with the algorithmic issue of solving (1.7) (or a penalized version thereof) in the original $p$-dimensional space. As pointed out by Wu et al. (2015), the latter leads to meaningless solutions with the objective function equal to infinity. Finally, we remark that the requirement in (2.17) is easy to satisfy as ${\rm rank}({\mathbb{X}}\widehat{B})\leq L$ and ${\rm rank}(\mathbb{I}_{n}-P_{{\mathbb{Y}}})=n-L$ when we observe at least one data point per class.
@@ 3 Theoretical guarantees
In this section we provide theoretical guarantees for our procedure. We start by listing the main assumptions used in our analysis. The following condition on the priors $\pi=(\pi_{1},\ldots,\pi_{L})$ is assumed.
@@
This is a common regularity assumption even in binary classification problems. We may relax Assumption 1 to $\min_{k\in[L]}\pi_{k}\geq c\log(n)/n$ in our analysis (see, for instance, Lemma 15 of Appendix C) at the cost of more technical proofs and a less transparent presentation.

### v2

@@ 2 Methodology
This, in turn, requires to estimate the discriminant directional matrix $B^{*}=\Sigma_{w}^{-1}{M}$. The following lemma establishes the connection between $B^{*}$ and the regression matrix $B$ in (1.12). Recall the definitions of $\pi$, $D_{\pi}$ and ${M}$ from (1.3) as well as $\Sigma$ from (1.4).
@@
Under model (1.1), we have
@@
$\Sigma_{XY}=\mathbb{E}(XY^{\top})-\mathbb{E}(X)(\mathbb{E}(Y))^{\top}={M}D_{\pi}$ (2.4)
@@
$\Sigma=\Sigma_{w}+\Sigma_{b}=\Sigma_{w}+{M}D_{\pi}{M}^{\top},$ (2.6)
@@
$\Sigma_{w}^{-1}=\Sigma^{-1}+\Sigma^{-1}{M}\left(\mathbf{I}_{L}-D_{\pi}{M}^{\top}\Sigma^{-1}{M}\right)^{-1}D_{\pi}{M}^{\top}\Sigma^{-1}.$
@@
$B^{*}$ $=\Sigma_{w}^{-1}{M}$
$=\Sigma^{-1}{M}+\Sigma^{-1}{M}\left(\mathbf{I}_{L}-D_{\pi}{M}^{\top}\Sigma^{-1}{M}\right)^{-1}D_{\pi}{M}^{\top}\Sigma^{-1}{M}$
$=\Sigma^{-1}{M}\left(\mathbf{I}_{L}-D_{\pi}{M}^{\top}\Sigma^{-1}{M}\right)^{-1}$
$=\Sigma^{-1}{M}D_{\pi}\left(D_{\pi}-D_{\pi}{M}^{\top}\Sigma^{-1}{M}D_{\pi}\right)^{-1}$
$=B\left(D_{\pi}-D_{\pi}{M}^{\top}B\right)^{-1}$
@@ 2.1 Estimation of the regression matrix
To accommodate high-dimensional data, we estimate $B$ via regularized regression methods in (1.14). Choices of the penalty term $\text{pen}(B)$ depend on the concrete problem at hand. For instance, when entries of $B$ exhibit certain smoothness structure such as in some imaging or audio applications (Hastie et al., 1995), one could choose $\text{pen}(B)={\rm tr}(B^{\top}\Omega B)$ for some pre-specified “roughness” penalty matrix $\Omega$. In this paper we mainly focus on the following two types of structural estimation of $B$.
Sparsity-based structural estimation.
If we believe that only a subset of the original features are important for predicting the label, then sparsity structures of the regression matrix $B$ are reasonable assumptions. If the subset of features, that are pertinent for predicting the label, varies across different levels, then the entry-wise lasso penalty is appropriate. The corresponding matrix estimator $\widehat{B}$ is defined in (1.14) with the following lasso penalty (Tibshirani, 1996)
@@
for some tuning parameter $\lambda>0$. In Section 3.3 we provide a complete analysis of this estimator. If there exist many features that are not predictive for all levels of the response label, then the group-lasso penalty $\text{pen}(B)=\lambda\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ (Yuan and Lin, 2006) is a more suitable choice. In this case, one can retain a much smaller subset of important features and achieve feature selection. Furthermore, one can also consider either a row-wise or a column-wise elastic net penalty (Zou and Hastie, 2005) if the features tend to (highly) correlate with each other. Our numerical results in Sections 4 and 5 include both the lasso penalty and the elastic net penalty.
Low-rank structural estimation.
Different from sparsity, if the goal is to construct a low-dimensional subset of linear combinations of the original feature for classification purpose, then a low-rank penalty in (1.14) could be more appropriate. For example, if the matrix of conditional means ${M}\in\mathbb{R}^{p\times L}$ has a reduced-rank $r\ll\min\{p,L\}$, then Lemma 1 implies that ${\rm rank}(B)\leq r$. In this case estimation of $B\in\mathbb{R}^{p\times L}$ in (1.14) could be done by using the following reduced-rank penalty (see, for instance, Izenman (1975))
$\text{pen}(B)=\lambda~{}{\rm rank}(B)$ (2.8)
for some tuning parameter $\lambda>0$. A full theoretical analysis of using (2.8) is presented in Section 3.4. Some alternative penalties include, but are not limited to, the reduced-rank penalty plus the matrix ridge penalty in Mukherjee and Zhu (2011), the nuclear norm penalty in Koltchinskii et al. (2011), adaptive nuclear norm penalty in Chen et al. (2013) as well as the generalized cross-validation rank penalty in Bing and Wegkamp (2019).
Remark 1.
We conclude this section by giving another concrete example under which $B$ has low rank. In Bing and Wegkamp (2023) the authors consider the model $X=AZ+W$ with latent factors $Z\in\mathbb{R}^{r}$ for some $r\ll p$ and some general additive noise $W$, that is independent of both $Z$ and the label $Y$. Conditioning on $Y={\bm{e}}_{\ell}$ for all $\ell\in[L]$, the low-dimensional factors $Z$ are assumed to follow Gaussian distributions with different conditional means $\alpha_{\ell}\in\mathbb{R}^{r}$ but the same within-class covariance matrix $\Sigma_{z}\in\mathbb{R}^{r\times r}$. When $W$ is also Gaussian, this model is a sub-model of (1.1) with $\mu_{\ell}=A\alpha_{\ell}$ and $\Sigma_{w}=A\Sigma_{z}A^{\top}+{\rm Cov}(W)$. When $r\leq\min\{p,L\}$, it is easy to verify that ${\rm rank}(B)\leq r$. The two-step approach in Bing and Wegkamp (2023) uses the unsupervised Principal Component Analysis (PCA) to reduce the feature dimension and then base classification on the reduced dimension. It is entirely different from our approach in Section 2.2 but connected to the one in Section 2.3 in the way that our approach in Section 2.3 is also a two-step procedure but reduces the feature space in the first step in a supervised way via the regression step in (1.14). Numerical comparison of both procedures are also included in our simulation study in Appendix A.
@@ 2.2 Estimation of the discriminant matrix and the Bayes rule
The discriminant functions in (2.1) that form the basis of the Bayes classification rule (1.2) necessitate that we estimate the matrix of the conditional means ${M}$, the vector of prior probabilities $\pi$ as well as the discriminant directional matrix $B^{*}$. We adopt the following standard estimators
$\widehat{M}={\mathbf{X}}^{\top}{\mathbf{Y}}({\mathbf{Y}}^{\top}{\mathbf{Y}})^{-1},\qquad D_{\widehat{\pi}}={\rm diag}(\widehat{\pi})={1\over n}{\mathbf{Y}}^{\top}{\mathbf{Y}}$ (2.9)
to estimate ${M}$ and $D_{\pi}$, respectively. Recall from Lemma 1 that $B^{*}=BH^{-1}$. Since the estimator $\widehat{B}$ of $B$ is computed in (1.14), in view of the closed-form of $H$ in (2.3), we propose to estimate $H$ by the plug-in estimator
$\widehat{H}=D_{\widehat{\pi}}-{1\over n}\widehat{B}^{\top}{\mathbf{X}}^{\top}{\mathbf{X}}\widehat{B}={1\over n}\left({\mathbf{Y}}^{\top}{\mathbf{Y}}-\widehat{B}^{\top}{\mathbf{X}}^{\top}{\mathbf{X}}\widehat{B}\right).$ (2.10)
Note that ${\mathbf{X}}\widehat{B}$ is simply the in-sample fit of the regression in (1.14). Computation of ${\mathbf{X}}\widehat{B}$ thus can be done efficiently even in non-parametric regression settings, for instance, when the columns of ${\mathbf{X}}$ consist of basis expansions. Now, using $B^{*}=BH^{-1}$, we can further estimate $B^{*}$ by
@@
Here for any matrix $M$, $M^{+}$ denotes its moore-penrose inverse. Finally, in view of (2.1), we estimate the discriminant functions by
$\widehat{G}_{\ell}(x)=\widehat{\mu}_{\ell}^{\top}\widehat{B}^{*}_{\ell}-2x^{\top}\widehat{B}^{*}_{\ell}-2\log(\widehat{\pi}_{\ell}),\qquad\text{for all $\ell\in[L]$. }$ (2.12)
@@
Theoretically, we show in Theorem 5 of Section 3.2 that $\widehat{H}$ is invertible with overwhelming probability under the same mild condition that is required for consistent classification. In Theorem 6 of Section 3.3 and Theorem 8 of Section 3.4, the condition is verified for the lasso estimator, and the reduced rank estimator, of $B$, respectively. Moreover, it is worth mentioning that the inverse of $\widehat{H}$ always exists when any $\ell_{2}$-norm related penalty is deployed in (1.14).
@@ 2.3 Classification based on reduced dimensions
We start with the fact (its proof is stated in Section B.1 for completeness) that the Bayes rule in (1.2) is equivalent to
@@
$C_{b}:=B^{\top}{M}D_{\pi}{M}^{\top}B.$
For a given reduced dimension $1\leq K<L$, we can find $K$ discriminant directions in the space of $Z$ by solving, for each $k\in[K]$, the optimization problem
$\begin{split}\alpha_{k}=\operatorname*{arg\,max}_{\alpha\in\mathbb{R}^{L}}&~{}\alpha^{\top}C_{b}~{}\alpha\quad\text{ s.t.}~{}\alpha^{\top}C_{w}~{}\alpha=1,~{}\alpha^{\top}C_{w}~{}\alpha_{i}=0,~{}~{}\forall~{}i<k.\end{split}$
@@
$\widehat{C}_{b}=\widehat{B}^{\top}\widehat{M}D_{\widehat{\pi}}\widehat{M}^{\top}\widehat{B}={1\over n}\widehat{B}^{\top}{\mathbf{X}}^{\top}P_{{\mathbf{Y}}}{\mathbf{X}}\widehat{B},\qquad\widehat{C}_{w}={1\over n}\widehat{B}^{\top}{\mathbf{X}}^{\top}(\mathbf{I}_{n}-P_{{\mathbf{Y}}}){\mathbf{X}}\widehat{B},$
with $P_{{\mathbf{Y}}}={\mathbf{Y}}({\mathbf{Y}}^{\top}{\mathbf{Y}})^{-1}{\mathbf{Y}}^{\top}$. Subsequently, the estimated discriminant directions $\widehat{\alpha}_{k}$, for $1\leq k\leq K$, are computed by solving
$\begin{split}\widehat{\alpha}_{k}=\operatorname*{arg\,max}_{\alpha\in\mathbb{R}^{L}}&~{}\alpha^{\top}\widehat{C}_{b}~{}\alpha\quad\text{ s.t.}~{}\alpha^{\top}\widehat{C}_{w}~{}\alpha=1,~{}\alpha^{\top}\widehat{C}_{w}~{}\widehat{\alpha}_{i}=0,~{}~{}\forall~{}i<k,\end{split}$ (2.17)
@@
$\text{span}({\mathbf{X}}\widehat{B})\subset\text{span}({\mathbf{X}})$
@@
${\rm rank}((\mathbf{I}_{n}-P_{{\mathbf{Y}}}){\mathbf{X}}\widehat{B})={\rm rank}({\mathbf{X}}\widehat{B}),$ (2.18)
we do not suffer numerically in solving (2.17), in stark contrast with the algorithmic issue of solving (1.6) (or a penalized version thereof) in the original $p$-dimensional space. As pointed out by Wu et al. (2015), the latter leads to meaningless solutions with the objective function equal to infinity. Finally, we remark that the requirement in (2.18) is easy to satisfy as ${\rm rank}({\mathbf{X}}\widehat{B})\leq L$ and ${\rm rank}(\mathbf{I}_{n}-P_{{\mathbf{Y}}})=n-L$ when we observe at least one data point per class.
@@ 3 Theoretical guarantees
In this section we provide a unified theory for analyzing the classifier $\widehat{g}$ in (2.13) based on estimating the regression coefficient matrix. In Section 3.1 we start with full generality and bound the excess risk of $\widehat{g}$ by a quantity that is related with the error of estimating $B^{*}$. The results there hold for any estimator $\widehat{B}^{*}$ of $B^{*}$. In Section 3.2 we focus on the proposed estimator $\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$ and provide a unified analysis of its estimation error that is valid for any generic estimator $\widehat{B}$. Finally we apply the general result to two particular choices of $\widehat{B}$, the lasso estimator in Section 3.3 and the reduced-rank estimator in Section 3.4.
3.1 A reduction scheme for bounding the excess risk
Let $(X,Y)$ be a new pair from (1.1) that is independent of the training data ${\bm{D}}:=\{{\mathbf{X}},{\mathbf{Y}}\}$. For any classifier $\widehat{g}:\mathbb{R}^{p}\to[L]$ in (2.13), its excess risk relative to the error rate of the Bayes rule in (1.2) is defined as
$\mathcal{R}(\widehat{g}):=\mathbb{P}\{Y\neq\widehat{g}(X)\mid{\bm{D}}\}-\mathbb{P}\{Y\neq g^{*}(X)\}.$
Note that we use the identification $\{Y=k\}:=\{Y={\bm{e}}_{k}\}$ in the above definition. Our analysis of $\mathcal{R}(\widehat{g})$ requires the following condition on the priors $\pi=(\pi_{1},\ldots,\pi_{L})$.
@@ 3.1 A reduction scheme for bounding the excess risk
This is a common regularity assumption even in binary classification problems. We may relax Assumption 1 to $\min_{k\in[L]}\pi_{k}\geq c\log(n)/n$ in our analysis (see, for instance, Lemma 21 of Appendix D) at the cost of more technical proofs and a less transparent presentation.
The following proposition states that for the purpose of bounding $\mathcal{R}(\widehat{g})$, it suffices to study the estimation error of each discriminant function, that is, $\widehat{G}_{\ell}(X)-G_{\ell}(X)$ for $\ell\in[L]$. The proof essentially follows the argument of proving Theorem 12 in Bing and Wegkamp (2023) but with modifications to capture the sum of squared errors of estimating all discriminant functions.
Proposition 2.
Under model (1.1) with 1, for any sequences $t_{1},\ldots,t_{L}$ (that possibly depend on ${\bm{D}}$), with probability equal to one, one has
$\mathcal{R}(\widehat{g})~{}\leq~{}C\sum_{\ell=1}^{L}t_{\ell}^{2}+L\sum_{\ell=1}^{L}\mathbb{P}\left\{|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\geq t_{\ell}\mid{\bm{D}}\right\}.$
Proof.
See Section B.2. ∎
2 is valid for any estimator $\widehat{G}_{\ell}$ of the discriminant function, in particular, for our proposed $\widehat{G}_{\ell}$ in (2.12). The choice of $(t_{1},\ldots,t_{L})$ depends on the estimation error of $|\widehat{G}_{\ell}(X)-G_{\ell}(X)|$. We therefore proceed to analyze
$\widehat{G}_{\ell}(X)-G_{\ell}(X)$ $=(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}+(\widehat{\mu}_{\ell}-2X)^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})+2\log\left({\pi_{\ell}/\widehat{\pi}_{\ell}}\right)$ (3.2)
for each $\ell\in[L]$. It is worth mentioning that in this section we allow $\widehat{B}^{*}$ in the above display to be any estimator of $B^{*}$ whereas $\widehat{\mu}_{\ell}$ and $\widehat{\pi}_{\ell}$ are given in (2.9). Since the variance of the first term in (3.2) depends on $\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$, we define
$\Delta_{\infty}:=\max_{\ell\in[L]}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}.$
This quantity essentially reflects the largest Mahalanobis distance in the feature space between any two label classes as can be seen from the relation (under $\mathbb{E}(X)=0_{p}$)
1, the Gaussian tail of $X\mid Y$ and the independence between $X$ and $({\mathbf{X}},{\mathbf{Y}})$ in (3.2) entail the following upper bound of $|\widehat{G}_{\ell}(X)-G_{\ell}(X)|$.
Theorem 3.
Under model (1.1) with 1, assume $\Delta_{\infty}\geq 1$ and $L\log(n)\leq n$. With probability at least $1-3n^{-1}$, we have, for all $\ell\in[L]$,
$\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|~{}\lesssim~{}\sqrt{\Delta_{\infty}}\sqrt{L\log(n)\over n}$ $+\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}+\log(n)}$
$+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|.$
Proof.
See Section B.3. ∎
Condition $L\log(n)\leq n$ together with 1 ensures that $\widehat{\pi}_{\ell}\asymp\pi_{\ell}$ for all $\ell\in[L]$. The reasonable condition $\Delta_{\infty}\geq 1$ requires that the largest separation among all pairs of conditional distributions does not vanish. It is assumed to simplify the presentation, and our analysis can be easily extended to the case $\Delta_{\infty}\to 0$ as $n\to\infty$.
Combining 2 with Theorem 3 immediately yields the following corollary. Denote by $\mathbb{E}_{{\bm{D}}}$ the expectation taken with respect to the training data ${\bm{D}}$ only. For any estimator $\widehat{B}^{*}$ of $B^{*}$, define the following (random) quantity
$\mathcal{Q}(\widehat{B}^{*}):=\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}^{2}(\Delta_{\infty}+\log n)+\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}.$ (3.3)
Corollary 4.
Under model (1.1) with 1, assume $\Delta_{\infty}\geq 1$. For any estimator $\widehat{B}^{*}$ of $B^{*}$, the corresponding classifier $\widehat{g}$ in (2.13) satisfies
$\mathbb{E}_{\bm{D}}\left[\mathcal{R}(\widehat{g})\right]~{}\lesssim~{}\mathbb{E}_{{\bm{D}}}\left[\min\left\{{\Delta_{\infty}}{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*}),~{}1\right\}\right].$
Proof.
See Section B.4. ∎
4 ensures that bounding $\mathcal{R}(\widehat{g})$ can be reduced to controlling $\mathcal{Q}(\widehat{B}^{*})$ for any estimator $\widehat{B}^{*}$ of $B^{*}$. In particular, 4 is valid for any existing approach based on estimating $B^{*}$. In the next section we provide a unified analysis of $\mathcal{Q}(\widehat{B}^{*})$ for our proposed estimator $\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$ based on a generic $\widehat{B}$. As shown later, the first term in (3.3) is typically the dominating term.
3.2 A unified analysis of the estimation of $B^{*}$ based on generic regression coefficient estimation
In this section we provide more explicit bounds of $\mathcal{Q}(\widehat{B}^{*})$ in (3.3) for the estimator $\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$ with $\widehat{H}$ obtained from (2.10). Our results in this section are valid for any estimator $\widehat{B}$ of $B$ that satisfies the following assumption.
Assumption 2.
There exists some deterministic sequences $\omega_{1}$ and $\omega_{2}$ such that the following holds with probability at least $1-n^{-1}$,
$\max\left\{n^{-1/2}\|{\mathbf{X}}(\widehat{B}-B)\|_{F},~{}\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}\right\}\leq\omega_{2},\quad\|\widehat{B}-B\|_{1,2}\leq\omega_{1}.$
We need the following regularity conditions on $\Sigma_{w}$. It assumes the largest entry of $\Sigma_{w}$ is bounded from above and the smallest eigenvalue of $\Sigma_{w}$ is bounded away from zero.
Assumption 3.
There exists some positive constants $0<c\leq C<\infty$ such that $c<\lambda_{p}(\Sigma_{w})\leq\|\Sigma_{w}\|_{\infty}\leq C$.
Theorem 5.
Under model (1.1) with Assumptions 1 & 3, assume $\Delta\asymp 1$ and $L\log(n)\leq cn$ for some small constant $c>0$. For any $\widehat{B}$ satisfying 2 with $\omega_{2}\sqrt{L}\leq c,$ with probability at least $1-2n^{-1}$,
the matrix $\widehat{H}$ in (2.10) is invertible,
and
$\mathcal{Q}(\widehat{B}^{*})\lesssim~{}L^{2}{\log^{2}(n)\over n}+\omega_{2}^{2}L^{2}\log(n)+\omega_{1}^{2}L^{3}{\log(n\vee p)\over n}.$ (3.4)
Proof.
See Section B.5. ∎
Theorem 5 relates $\mathcal{Q}(\widehat{B}^{*})$ to the estimation error of $\widehat{B}$ in both the Frobenius norm and the $\ell_{1}/\ell_{2}$-norm. In the bound of (3.4), the second term on the right-hand-side is usually the leading term and its rate depends on specific choice of $\widehat{B}$. We provide concrete examples in Section 3.3 for the lasso estimator and in Section 3.4 for the reduced-rank estimator.
Condition $\omega_{2}\sqrt{L}\leq c$ ensures that, with high probability, $\lambda_{K}(\widehat{H})\geq\lambda_{K}(H)/2$ holds so that $\widehat{H}$ is invertible. Furthermore, the bound in (3.4) suggests that the condition $\omega_{2}\sqrt{L}\leq c$ is also needed for consistent classification.
3.3 Application to estimation of $B$ with entry-wise $\ell_{1}$-regularization